# Line Search Strategy for Navigating through Barren Plateaus in Quantum Circuit Training

### v2

@@ 5 Results
In this section, we present the outcomes of our numerical simulations conducted on variational problems, specifically the estimation of the ground state of an XXX Heisenberg Hamiltonian and the SYC model. In both of these models, the ground state satisfies volume law, indicating that the entanglement entropy of any bipartition of the state scales with the volume, i.e., $(\rho_{A})\sim|A|$ (refer to Ref. [39] for a comprehensive review of these concepts). Additionally, we share numerical results obtained from quantum gate decomposition problem scenarios. Our findings demonstrate that the developed evolutionary algorithm excels in deep quantum circuit synthesis tasks, outperforming traditional optimization methods such as the well-known Broyden–Fletcher–Goldfarb–Shanno (BFGS) [62] or the ADAM [61] optimization methods. These traditional methods prove ineffective in producing meaningful results for the gate decomposition.
@@
Our initial objective in these numerical experiments is to demonstrate the efficient utilization of the gradient-free multi-parameter search direction in circumventing BPs during circuit training. To achieve this, we focus on a specific scenario in which only a single agent operates on a randomly chosen subset of the parameters. In this form, the strategy closely resembles traditional gradient descend optimization algorithms. The key difference is the incorporated routine to determine the search direction. While conventional gradient-based optimizers choose the search direction based on local properties of the cost-function landscape, such as the gradient or curvature, our proposed alternative approach suggests that determining the search direction via long-range properties of the optimization landscape yields superior results, without introducing extra hyper-parameters in the optimization strategy. Our second objective was to enhance the optimization strategy by introducing mechanisms to prevent or significantly reduce the likelihood of getting trapped in local minima within the optimization landscape. To achieve this, we employed a multi-agent approach and implemented evolutionary selection among them.
@@
Figure 5: VQE simulations implementing the strategy based on the gradient-free multi-parameter search direction with a single agent. The training of the quantum circuits was conducted to find the ground state for the $N=10$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of sub-figure (a) and for a $10$-qubit SYC model. Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy evaluated on the sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model. Sub-figures (c) and (d) show the same kind of results for the SYC model. The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-21.28$ eV and $S_{2}/S_{Page}=0.667$. The exact ground state of the SYC model is $E_{0}=-3.82$ eV with normalized entropy $S_{2}/S_{Page}=0.980$. For the initialization we used random parameters.
@@
From theoretical point of view, the Sachdev–Ye–Kitaev (SYK) model- [48, 49, 50], with it’s non-local Hamiltonian structure in the Puali basis, is a frequently visited representative of variational problems that are proven to be hard to solve in VQE context. This is because the SYC model features a ground state that is nearly maximally entangled [66, 50], so crossing a BP becomes inevitable during the VQE process [36]. From solid state physics point of view, the SYK model can be used to describe $2N$ spinless Majorana fermions $\chi_{i}$ satisfying the following anticommutation relations $\{\chi_{i},\chi_{j}\}=\delta_{i,j}$. The Hamiltonian of the model reads as
@@
In Fig. 5, we present the results of our numerical simulations conducted on $10$-qubit Heisenberg XXX and SYC models. These systems have been recognized as challenging for VQE in numerous studies, and our tests with conventional gradient-based optimizers affirmed these observations. To address the optimization problem (2), we explored four strategies available within the SQUANDER package. The Powell-variant of the BFGS optimizer [67] and the Gradient Descent methods incorporate an inexact line search [40] in each iteration, proving to be quite efficient for smaller problems (e.g., in the synthesis of $3-5$ qubit circuits). The handcrafted implementations of these optimizers also consider the periodicity of the optimization landscape by appropriately bounding the line search during optimization. Additionally, we included the ADAM optimization algorithm in our tests, although we did not implement the adaptive learning rate strategy proposed in [39]. Our findings indicate that, despite exhaustive optimization and adaptation to quantum circuit training circumstances, these optimization engines struggle to solve these optimization problems at scale.
In contrary, our developed optimization algorithm demonstrated the ability to mitigate the effects of BPs in situations where other optimization methods fail. Notably, our approach avoids BPs without the need for external interventions or monitoring of specific quantities. As anticipated, in our numerical simulations, gradient-based optimization approaches became trapped near the energy level of the initial point.
The left side of Fig.5 illustrates our VQE experiments on a $10$-qubit Heisenberg model identical to the system studied in [39]. In each iteration step, we performed optimization by randomly selecting $64$ parameters of the circuit and conducted a line search along the direction determined by Eq.(9). Throughout the iterations, we utilized a single agent, and no evolutionary steps were applied during the simulations. Still, the optimization problem could be solved, while significantly reducing the computational complexity.
@@
Regarding the exact solution of the SYC model, we need to mention that the diagonalization of the SYC Hamiltonian revealed a $2$-fold degeneracy of the states. In addition, the $2$-fold ground state and the nearest excited states (also $2$-fold degenerate) were close to each other in energy. While the energy of the ground state was $E_{0,SYC}=-3.819$ eV, the first excited energy was $E_{1,SYC}=-3.799$ eV. The VQE energy obtained with the training of the $p=100$ layered circuit turned out to be $E_{VQE}=3.814$ eV, a value that falls between the ground state energy and the first excited energy. Thus, the VQE state is expected to have a finite component orthogonal to the subspace of the $2$-fold ground state. Fortunately, this orthogonal component has only an overlap $M=0.098$ with the subspace of the first excited states, while the overlap of the VQE wave function with the subspace of the $2$-fold ground state is $M=0.902$. Hence, the components of the VQE wave function predominantly come from the ground state and the first excited state.
At this point we would like to also mention that the optimization Ref. [39] was not able to solve neither the same Heisenberg model nor the SYC model of identical size with an accuracy comparable to our results. (We also notice that [39] used slightly different circuit ansatz compared to ours.)
@@
Figure 6: VQE simulations implementing the strategy based on the gradient-free multi-parameter search direction with a single agent. The training of the quantum circuits was conducted to find the ground state for the $N=16$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of subfigure (c). Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy on sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model evaluated with shallower circuits consisting of $p=50$ and $100$ layers, while . Subfigures (c) and (d) show the same kind of results obtained with deep circuits consisting of $p=500$ and $1000$ layers. The black numbers and dots represent the evolution of the of the overlap integral between the VQE wave function and the exact ground state defined by Eq.(17). The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-32.84$ eV and $S_{2}/S_{Page}=0.693$. The optimization was initialized with parameters equal to 00.
@@
After examining $10$-qubit systems, we raise the stakes and turn our attention to even more challenging problems. While it is not a significant challenge for quantum computer simulators to execute circuits beyond $20$ qubits, the VQE experiments become quite expensive at this scale due to the large number of cost function evaluations. Therefore, in this work, we limit our considerations to problems up to $16$ qubits. The goal of our VQE experiments is to reproduce the ground state of a randomly generated $16$-qubit XXX Heisenberg model (with $k=3$), as visualized in the inset of Fig. 6.(c). Furthermore, to increase the efficiency of the optimization process, we initiate the training of the circuits with parameter values equal to 00. Since the circuits are applied to the initial state $|\mathbf{0}\rangle$ and initially, all the single-qubit rotations will be identities with this choice, the two-qubit entangling gates will lose their effect as well. This is because all the control qubits are initially in state 00, so the controlled two-qubit gates incorporated into the circuit will act as an identity gate as well. Consequently, the resulting state will have zero initial entanglement entropy, promising a steep initial descent in the VQE energy. Despite of this simplification, our method remains autonomous in avoiding BPs on its own, there is still no need for any external intervention during the optimization.
@@
The right panels of Fig.6 show our numerical results obtained with these deep circuits composed of $p=500$ and $p=1000$ layers. As the plotted results indicate, the developed optimization strategy had no difficulties proceeding with the training of these deep circuits. We still executed only a single agent in these numerical experiments, and optimized $128$ parameters in each iteration. We observed that the efficiency of the optimization varies from run to run. Fig.6 shows the most successful executions, while in several experiments, we observed signs of convergence to a local minimum around the energy level of $\sim-29$ eV. We argue that this issue could be resolved by increasing the number of agents in the process due to the evolutionary selection of the most successful agents. However, due to the overwhelming numerical requirements of the performed VQE simulations, we did not try out this approach in these simulations. We revisit this aspect of the algorithm in the next section by addressing quantum gate synthesis problems, where it becomes essential to avoid local minima to retrieve meaningful results. In these experiments, evolutionary selection will play a crucial role in solving the addressed problems.
@@
It is worth to study the computational complexity of the developed evolutionary algorithm described in this study. There are two key components in relation with our optimization technique characterizing it’s cost. These are the utilization of the agents undergoing evolutionary selections and the size of the parameter subset over which the cost function gets optimized. Both of these dimensions, though, fall back to the count of sampling from the single parameter cross sections of the cost function. Since the terminating point of the line search can be reused in the next iteration, only $2$ cost function evaluations are needed to reconstruct the parameter dependence of the cost function, if Eq. (5) is used to characterize the optimization landscape. In the more general case of Eq. (7) $4$ cost function evaluations are required to determine the minimum with respect to a single parameter. Therefore, $2LM$ (or $4LM$) cost function evaluations are required to determine the search direction, and additional constant number of evaluations to perform the (inexact) line search. These numbers compare close to the case of regular gradient descend method. In particular, on a quantum computer the general rule of parameter shift [75] can be used to evaluate a gradient component of the cost function with $2$ evaluations of the cost function. By carefully planning the number or agents and the number of optimized parameter components per iteration, the overall complexity of our optimization algorithm can be set not to exceed the complexity of conventional gradient based methods.

### v3

@@ 5 Results
In this section, we present the outcomes of our numerical simulations conducted on variational problems, specifically the estimation of the ground state of an XXX Heisenberg Hamiltonian and the SYC model. In both of these models, the ground state satisfies volume law, indicating that the entanglement entropy of any bi-partition of the state scales with the volume, i.e., $(\rho_{A})\sim|A|$ (refer to Ref. [39] for a comprehensive review of these concepts). Additionally, we share numerical results obtained from quantum gate decomposition problem scenarios. Our findings demonstrate that the developed evolutionary algorithm excels in deep quantum circuit synthesis tasks, outperforming traditional optimization methods such as the well-known Broyden–Fletcher–Goldfarb–Shanno (BFGS) [62] or the ADAM [61] optimization methods. These traditional methods prove ineffective in producing meaningful results for the gate decomposition.
@@
Our initial objective in these numerical experiments is to demonstrate the efficient utilization of the gradient-free batched search direction in circumventing BPs during circuit training. To achieve this, we focus on a specific scenario in which only a single agent operates on a randomly chosen subset of the parameters. In this form, the strategy closely resembles traditional gradient descent optimization algorithms. The key difference is the incorporated routine to determine the search direction. While conventional gradient-based optimizers choose the search direction based on local properties of the cost-function landscape, such as the gradient or curvature, our proposed alternative approach suggests that determining the search direction via long-range properties of the optimization landscape yields superior results, without introducing extra hyper-parameters in the optimization strategy. Our second objective was to enhance the optimization strategy by introducing mechanisms to prevent or significantly reduce the likelihood of getting trapped in local minima within the optimization landscape. To achieve this, we employed a multi-agent approach and implemented evolutionary selection among them.
@@
Figure 5: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=10$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of sub-figure (a) and for a $10$-qubit SYC model. Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy evaluated on the sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model. Sub-figures (c) and (d) show the same kind of results for the SYC model. The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-21.28$ eV and $S_{2}/S_{Page}=0.667$. The exact ground state of the SYC model is $E_{0}=-3.82$ eV with normalized entropy $S_{2}/S_{Page}=0.980$. For the initialization we used random parameters.
@@
From theoretical point of view, the Sachdev–Ye–Kitaev (SYK) model- [66, 67, 68], with it’s non-local Hamiltonian structure in the Puali basis, is a frequently visited representative of variational problems that are proven to be hard to solve in VQE context. This is because the SYC model features a ground state that is nearly maximally entangled [69, 68], so crossing a BP becomes inevitable during the VQE process [36]. From solid state physics point of view, the SYK model can be used to describe $2N$ spinless Majorana fermions $\chi_{i}$ satisfying the following anti-commutation relations $\{\chi_{i},\chi_{j}\}=\delta_{i,j}$. The Hamiltonian of the model reads as
@@
Firstly, in Fig. 5 we present the results of our numerical simulations conducted on a $10$-qubit Heisenberg XXX and SYC models. We wanted to address the same physical systems as in prior study of Ref. [39] giving us the opportunity to compare our method to what was described in [39]. The strategy developed in [39] work relies on the monitoring of the second-Rényi entropy (10) and reducing the learning rate when the entropy builds on high values.
Though, our numerical results revealed that the addressed system is still too small without the requirement of any special treatments to mitigate the BP effects, provided the VQE energy can be obtained with ideal measurement. Notably, our approach (and other gradient based strategies) is capable to solve the optimization problem without the need for external interventions or monitoring of specific quantities: during the calculations the second-Rényi entropy autonomously remained at low values.
The left side of Fig.5 illustrates our VQE experiments on a $10$-qubit Heisenberg model identical to the system studied in [39]. In each iteration step, we performed optimization by randomly selecting $L=64$ parameters of the circuit and conducted a line search along the direction determined by Eq.(9). Throughout the iterations, we utilized a single agent, and no evolutionary steps were applied during the simulations. Still, the optimization problem could be solved, while significantly reducing the computational complexity.
@@
Regarding the exact solution of the SYC model, we need to mention that the diagonalized SYC Hamiltonian revealed a $2$-fold degeneracy of the states. In addition, the $2$-fold ground state and the nearest excited states (also $2$-fold degenerate) were close to each other in energy. While the energy of the ground state was $E_{0,SYC}=-3.819$ eV, the first excited energy was $E_{1,SYC}=-3.799$ eV. The VQE energy obtained with the training of the $p=100$ layered circuit turned out to be $E_{VQE}=-3.814$ eV, a value that falls between the ground state energy and the first excited energy. Thus, the VQE state is expected to have a finite component orthogonal to the subspace of the $2$-fold ground state. Fortunately, this orthogonal component has only an overlap $M=0.098$ with the subspace of the first excited states, while the overlap of the VQE wave function with the subspace of the $2$-fold ground state is $M=0.902$. Hence, the components of the VQE wave function predominantly come from the ground state and the first excited state.
@@
Figure 6: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=16$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of subfigure (c). Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy on sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model evaluated with shallower circuits consisting of $p=50$ and $100$ layers, while . Subfigures (c) and (d) show the same kind of results obtained with deep circuits consisting of $p=500$ and $1000$ layers. The black numbers and dots represent the evolution of the of the overlap integral between the VQE wave function and the exact ground state defined by Eq.(17). The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-32.84$ eV and $S_{2}/S_{Page}=0.693$. The optimization was initialized with parameters equal to 00.
@@
After examining $10$-qubit systems, we raise the stakes and turn our attention to even more challenging problems. Although it is not a significant challenge for quantum computer simulators to execute circuits beyond $20$ qubits, VQE experiments become quite expensive at this scale due to the large number of cost function evaluations. Therefore, in this work, we limit our considerations to problems up to $16$ qubits. The goal of our VQE experiments is to reproduce the ground state of a randomly generated $16$-qubit XXX Heisenberg model (with $k=3$), as visualized in the inset of Fig. 6.(c). Furthermore, to increase the efficiency of the optimization process, we initiate the training of the circuits with parameter values equal to 00. Since the circuits are applied to the initial state $|\mathbf{0}\rangle$ and initially, all single-qubit rotations will be identities with this choice, the two-qubit entangling gates will lose their effect as well. This is because all control qubits are initially in state 00, so the controlled two-qubit gates incorporated into the circuit will act as an identity gate as well. Consequently, the resulting state will have zero initial entanglement entropy, promising a steep initial descent in the VQE energy. Despite this simplification, our method remains autonomous in avoiding BPs on its own, there is still no need for any external intervention during the optimization.
@@
The right panels of Fig.6 show our numerical results obtained with these deep circuits composed of $p=500$ and $p=1000$ layers. As the plotted results indicate, the developed optimization strategy had no difficulties proceeding with the training of these deep circuits. We still executed only a single agent in these numerical experiments, and optimized $128$ parameters in each iteration. We observed that the efficiency of the optimization varies from run to run. Fig.6 shows the most successful executions, while in several experiments, we observed signs of convergence to a local minimum around the energy level of $\sim-29$ eV. We argue that this issue could be resolved by increasing the number of agents in the process due to the evolutionary selection of the most successful agents. However, due to the overwhelming numerical requirements of the performed VQE simulations, we did not try out this approach in these simulations. We revisit this aspect of the algorithm in a later section by addressing quantum gate synthesis problems, where it becomes essential to avoid local minima to retrieve meaningful results. In these experiments, evolutionary selection will play a crucial role in solving the addressed problems.
@@
Comparison to conventional gradient descent and gradient free methods:
To assess the computational efficiency of our novel optimization method, we conduct a comparative analysis with traditional gradient descent (GD) approaches. We begin by evaluating the optimization capabilities of each method on a smaller system, and then scale up to a larger system to examine how their efficiency holds up under increased complexity. On the lower end of the range we refer to our results plotted in Fig. 7.
Figure 7: Performance comparison of $4$ optimization strategies to solve Eq.(2) on a $10$-qubit Heisenberg model identical to the system plotted in Fig.5. The blue colored line stands for the optimization method developed by this work with $L=64$. Red lines correspond to basic gradient descent optimization with a learning rate $\eta=0.001$. Orange lines represent results obtained with a gradient descent method (GS) with inexact line search (ILS) and the same learning rate parameter. Brown colored lines show the results obtained with the ADAM optimization method with a learning rate $\eta$.
To address the optimization problem (2), we explored four strategies available within the SQUANDER package. The Powell-variant of the BFGS optimizer [70] and the GD methods incorporate an inexact line search[40] in each iteration, proving to be quite efficient for smaller problems (e.g., in the synthesis of $3-5$ qubit circuits). Though, the BFGS method, as a second order method is loosing the performance at large parameter count, hence we omit this method from our comparison on VQE experiments. The handcrafted implementations of these optimizers also consider the periodicity of the optimization landscape by appropriately bounding the line search during optimization. Additionally, we included the ADAM optimization algorithm in our tests, although we did not implement the adaptive learning rate strategy proposed in [39]. In Fig.7. we see that the examined optimization engines has similar performance in solving the optimization problem, with the best performance attributed to the ADAM method and to the strategy developed by this work. These two optimizers had much bigger steepness in the cost function curve for both zero and random initializations. The dominant part of the optimization happened in the first quarter of the performed cost function evaluations, while further refinement of the results got significantly slowed down. The steep decrement in the cost function at the beginning can be translated into a reduction in the number of required cost function evaluations to achieve reasonable result, possibly make faster the optimization over BPs.
Figure 8: Performance comparison of $4$ gradient free optimization strategies to solve Eq.(2) on a $10$-qubit Heisenberg model identical to the system plotted in Fig.5. The blue colored line stands for the optimization method developed by this work with $L=64$. Red lines correspond to the Cobyla method, orange lines represent results obtained with the Nelder-Mead optimization and the brown line stands for the results obtained with the Powel gradient free method. The gradient free optimizers are taken fro the SciPy package.
As our newly developd method reminds the basic concepts of derivative-free approaches, we also explore the capabilities of gradient-free algorithms for comparison. (We took the implementations of the explored gradient-free optimization methods from the SciPy library.) However, we don’t consider our method as a gradient-free approach because our method reproduces the whole cross section of the optimization landscape along single parameter directions. The information on the gradient components also can be accessed with no further computation costs.
In Fig. 8. we show the results of our numerical simulations on the same $10$-qubit system as before. We found that two gradient-free optimizers, Nelder-Mead[71] and Cobyla[72], were unable to solve the $10$-qubit optimization problem. This is in contrast to a prior study[18] reporting adequate capability for these optimizers in solving similar problems. However, there are key differences between the two studies. For example, Ref. [18] used a limited number of parametric layers, whereas our study used a larger number of layers in the circuit ansatz. Secondly, the objective of Ref.[18] was to reproduce the identity matrix from a circuit instantiated with randomly generated parameters. Thus, the target of the optimization in Ref. [18] is presumably not located on a BP either.
In contrast, Powell’s gradient free optimization method[73] showed surprisingly good efficiency in training the addressed quantum circuits. In fact, Powell’s method shares many similarities with our approach, specifically in its manner of use of line searches. The method begins by defining an initial set of search directions, typically coordinate directions in the parameter space. It then iteratively searches along each of these directions, seeking the optimal displacement that minimizes the cost function while only adjusting the parameters along that specific direction. This process typically employs a variant of the line search algorithm to determine the optimal displacements. After exploring all the search directions, the method concludes the iteration by replacing the search direction corresponding to the largest displacement with a linear combination of the original search directions. This process is repeated iteratively during the optimization, with the iteration terminating when a convergence criterion is met.
However, our method has some advantages over Powell’s method. For example, our method can determine the direction of the search direction in parallel, which allows for concurrent executions and reduces the execution time. Additionally, our method takes advantage of the properties of the optimization landscape to determine the maximal distance of the line search.
Powell’s method was able to solve the optimization problem efficiently, but it exhibited abrupt oscillations in the second-Rényi entropy, which may indicate a loss of efficiency on larger problems. Since the addressed optimization problem is relatively small, it’s too early to draw conclusions about the efficiency of the best-performing optimization strategies at scale. We tried to fill up this gap via further numerical experiments performed on larger systems. Our numerical results showed that by increasing the system size the efficiency of most of the optimization engines start to drop radically. We showcase this tendency via numerical experiments performed on the $16$-qubit Heisenberg model depicted in Fig.6 with $p=500$ circuit layers. This model delivers high degree of computational complexity for a well established numerical investigations.
Figure 9: Performance comparison of $4$ optimization strategies to solve Eq.(2) on a $16$-qubit Heisenberg model identical to the system plotted in Fig.5. The blue colored line stands for the optimization method developed by this work with $L=64$. Red lines correspond to basic gradient descent optimization with a learning rate $\eta=0.0001$. Orange lines represent results obtained with a gradient descent method (GS) with inexact line search (ILS) and the same learning rate parameter. Brown colored lines show the results obtained with the ADAM optimization method with a learning rate $\eta$. Finally, green color stands for the results of Powell’s gradient-free optimizer.
Our analysis of the larger system reveals a significant difference in optimization performance when starting from random versus (near-)zero initial parameter vectors. When initialized randomly, all five examined optimizers struggled to solve the optimization problem [as shown in Fig.9.a)]. This is because the optimizations began at a BP, making it difficult to approach the global minimum of the VQE energy. The second-Rényi entropy remained almost constant throughout the iterations [Fig.9.c)], indicating that the trajectories did not escape the BP. As a result, the VQE energies improved only marginally as the number of cost function evaluations increased. Notably, the method developed in this work still demonstrated the best performance in solving the optimization problem.
When initializing the optimization with identity operators (i.e., starting with zero rotational parameters), the second-Rényi entropy begins to evolve from a small value, potentially enabling a steeper decrease in the cost function [as shown in Fig. 9.d)]. Still, our comparison reveals that gradient-based optimizers exhibit significantly lower efficiency in solving the $16$-qubit problem compared to the $10$-qubit VQE problem. This inefficiency is likely linked to the second-Rényi entropy, as the ADAM and GD optimizers, which build up small values of the second-Rényi entropy, are the most efficient among the gradient-based optimizers. In contrast, the gradient descent method with inexact line search enters a region of high entropy early in the optimization, making it highly inefficient. In contrast, the optimization method developed in this work maintains its efficiency observed for the $10$-qubit system during the first quarter of the iterations, rapidly approaching the target value of the optimization [as shown by the blue line in Fig. 9.b)]. However, when the second-Rényi entropy builds up higher values, the steepness of the VQE energy significantly reduces, making only a marginal improvement on the VQE energy. Unfortunately, our numerical experiments were unable to fully recover the target value of the exact ground state. As discussed earlier, the used circuit ansatz may not be optimized for the addressed physical system, potentially lacking the necessary expressiveness to capture the ground state. We leave the study of possible impacts of various circuit structures on the quality of the converged solution for future work. Powell’s method also showed outstanding efficiency in decreasing the cost function, it performed almost equally well as our optimization method. Nevertheless, Powell’s method has a significant limitation. Due to its sequential nature, it becomes impractical for large-scale problems, making it less suitable for VQE tasks at scale. Also, we should notice another limiting well pronounced steps (see the green lines in Fig. 9.a) and b)). In Powell’s method the search directions are determined iteratively from the search directions of the previous step and therefore the optimization along the search directions will affect more and more parameters at once as the optimization iterations go on. Due to our reasoning this is a key component in limiting the efficiency of an optimization method to solve the VQE problem at scale. To overcome this issue, we can restart the solver’s inner state and start over the optimization with search direction aligning with single parameter directions as we do in the optimization method developed in this work. The steps in the green curves in Fig. 9.a) and b) indicates the resets of Powell’s optimization method. This approach allows us to lift the limitation and continue the improvement on the cost function. The other two gradient free optimizers used in our $10$-qubit simulations did not performed well there, hence we omitted them from our larger experiments on $16$ qubits.
To understand the underlying factors contributing to the success of the optimization method presented in this work, we will analyze its performance and compare it to that of gradient-based optimizers. Specifically, we will investigate why our method outperforms the gradient-based optimizers in our experiments. To establish a baseline for comparison, we first examined the optimal learning rate parameter for the GD optimization algorithm.
Figure 10: Numerical experiments performed with the gradient descent algorithm with various learning rate parameters labeled by $\eta$. In each iterations all of the free parameters are updated.
In Fig.10, we visualize the evolution of the VQE energy as a function of cost function evaluation count. We observed pronounced oscillations in the cost function for larger learning rates, specifically for $\eta=10^{-2}$ and $\eta=10^{-3}$ (blue and red lines, respectively). Interestingly, the curve for $\eta=10^{-3}$ exhibits the larger oscillations. This phenomenon can be attributed to the optimization landscape’s narrow valleys, which are surrounded by flat territories, consistently with the barren plateau theory (i.e. the dominant part of the landscape is expected to be flat, with narrow valley that are hard to find). A larger learning rate is more likely to overstep these valleys, resulting in oscillations with smaller amplitudes. Conversely, a lower learning rate increases the likelihood of entering these valleys, but may also lead to larger steps during parameter updates (due to the increased gradient inside the valley), causing the algorithm to leave the valley. This reasoning explains the large oscillations in the red line. As the learning rate is further decreased, the oscillations are smoothed out, but the optimization performance is also reduced due to smaller updates in the parameter space (orange and brown lines). Our experiments suggest that learning rates between $\sim 10^{-4}$ and $\sim 10^{-3}$ are the most effective choice.
Figure 11: The efficiency of gradient descent (GD) based VQE optimization for various batch sizes $L$. The batch size refers to the number of parameters taken from the set of the free parameters to be updated during a single GD iteration. In the experiments we applied a learning rate of $\eta=0.001$. The algorithm reported in this work is indicated by the last entry in the legends.
Interestingly, the oscillations observed during optimization also disappear when updating only a smaller subset of the parameters in each optimization iteration. This means that, similarly to our gradient-free multi-parameter search direction approach developed in this work, we randomly select a finite subset of the parameter space in each iteration, evaluate the associated gradient components of the cost function, and apply the parameter update rule on the chosen subset. To illustrate this tendency, Fig.11 shows the evolution of the VQE energy for various batch sizes $L$. (The batch size refers to the number of the randomly chosen components in the parameters vector.) The circuit ansatz with $500$ layers has $45000$ free parameters. As we decrease the batch size from $45000$ to $500$, the oscillations in the cost function gradually disappear. While the high dimensionality of the optimization landscape makes it challenging to develop an exact explanation, we can offer a qualitative reasoning for this observation. By selecting a small subset of parameters to vary in each iteration, the circuit’s expressiveness gets narrowed down, effectively reducing the risk of getting onto a BP when initiated from zero parameters [20]. By optimizing a low expressive circuit in each iteration of the training process one can effectively avoid BPs during the optimization. However, this procedure does not fully solve the BP issue, as optimization over a selected subspace of the parameter space can only reach the minimum corresponding to that particular subspace. If already close to a local minimum (or on a BP), the improvement in the cost function will be small. These two competing aspects predestinates a batch size for which the optimization is the most efficient. Our numerical results in Fig.11 suggest that a batch size around $50$ is the most efficient choice for gradient descent optimization.
Now, we’d like to discuss the impact of parameter initialization on the optimization process. As discussed previously, when starting from a randomly initialized parameter set, the improvement from optimizing a randomly chosen parameter subspace is in general small. This is expected, as the starting point is likely to be located on a BP. In contrast, when starting the optimization with zero initialization (i.e., setting the single qubit rotation gates to identity operations), the VQE energy undergoes a much larger improvement during the first stage of the optimization. This is what we have exploited in our numerical results shown in Fig.11 as well. To avoid zero gradient components at the beginning of the optimization, we set the initial parameters to randomly selected, but small values close to zero. Our gradient-free optimization strategy can be started from exact zero initial parameters, allowing for even larger initial decrements in the VQE energy, as plotted by the purple lines of Fig.11. After a pronounced initial decrement in the cost function, the degree of further improvements in the VQE energy becomes modest, indicating the impact of BPs surrounding the global minimum of the optimization problem. The second-Rényi entropy of the exact ground state is relatively high [see Fig. 6.b)], suggesting the presence of a BP.
@@
It is worth to study the computational complexity of the developed evolutionary algorithm described in this study. There are two key components in relation with our optimization technique characterizing it’s cost. These are the utilization of the agents undergoing evolutionary selections and the size of the parameter subset over which the cost function gets optimized. Both of these dimensions, though, fall back to the count of sampling from the single parameter cross sections of the cost function. Since the terminating point of the line search can be reused in the next iteration, only $2$ cost function evaluations are needed to reconstruct the parameter dependence of the cost function, if Eq. (5) is used to characterize the optimization landscape. In the more general case of Eq. (7) $4$ cost function evaluations are required to determine the minimum with respect to a single parameter. Therefore, $2LM$ (or $4LM$) cost function evaluations are required to determine the search direction, and additional constant number of evaluations to perform the line search. These numbers compare close to the case of regular gradient descent method. In particular, on a quantum computer, the general rule of parameter shift [81] can be used to evaluate a gradient component of the cost function with $2$ evaluations of the cost function. By carefully planning the number of agents and the number of optimized parameter components per iteration, the overall complexity of our optimization algorithm can be set not to exceed the complexity of conventional gradient-based methods.